Exercise T4-1 Swifty WaterWorks manufactures snorkel gear. During the past month, Swifty purchased 4,200 pounds of plastic to use in its dive masks, at a cost of $7,140. The standard price for the plastic is $1.50 per pound, and each mask requires /4 pound of plastic. The company actually used 4,100 pounds of the plastic to produce 16,260 dive masks. (a) Prepare the journal entry to record the purchase of plastic during the month. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit To record the purchase of plastic) (b) Prepare the journal entry to record the use of plastic during the month. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit To record the use of plastic)
Answer #1
Debit Credit
a
Direct materials inventory 6300 =4200*1.5
Direct materials Price variance 840
        Accounts Payable 7140
b
Work in Process inventory 6097.50 =16260*1.5*1/4
Direct materials quantity variance 52.50
        Direct materials inventory 6150 =4100*1.5
Note: The amounts in entry b may alternatively be rounded off to nearest whole number.
6097.50 may be written as 6097 or 6098
52.50 may written as 52 or 53